High-Low Method

Shatner Inc. has decided to use the high-low method to estimate the total cost and the fixed and variable cost components of the total cost. The data for various levels of production are as follows:

Units Produced Total Costs
5,800 $694,400
1,800 446,400
3,720 673,070

a. Determine the variable cost per unit and the total fixed cost.

Variable cost: (Round to the nearest dollar.) $ per unit
Total fixed cost: $

b. Based on part (a), estimate the total cost for 2,740 units of production.

Total cost for 2,740 units: $
Answer #1

High - Low Method of Accounting.

a ) Variable cost per unit = ( $ 694400 - 446400 ) / ( 5800 - 1800 ) = $ 62.

At high volume

>> Fixed cost + ( 5800 * $ 62 ) = $ 694400.

>> Fixed cost = 334800.

b ) Total cost for 2740 units = 334800 + ( 2740 * 62 ) = $ 504680.